OpenStudy (anonymous):

What is the slope of the line? y = -3x – 4 A. -4 B. -3 C. 3 D. 4

OpenStudy (solomonzelman):

y=mx+b m is the slope. COMPARE!

OpenStudy (anonymous):

y = mx + b Where m is the slope. Just look at it.

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

m is a number in front of x, at least give it a try.

OpenStudy (solomonzelman):

Lets see, if you had y=7x-2, comparing to y=mx+b, 7x=mx m=7 WHAT about YOUR EXAMPLE?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

its -3....

OpenStudy (anonymous):

oh